Qx=150 -PxPo find point price and cross elasticities
1
Expert's answer
2019-04-09T08:49:54-0400

Qx=150 - PxPo, so Px = (150 - Qx)/Po

Point price elasticity is: Ed = -b*Px/Qx = -1/Po*Px/Qx = -Px/(Po*Qx).

To find cross-price elasticity of demand, we need to know changes in price Po and quantity Qx of two goods.